How many ways can you roll a die and flip a coin and end up with heads and an odd number?
3 ways, out of 12 possible outcomes.
How do you find the probability in math?
The formula for probability is as follows:
Probability = Total number of favorable cases / total number of cases
It always lies between 0 and 1 and has no unit.
The inclusion of a parity bit extends the message length. There are more bits that
can be in error since the parity bit is now included. The parity bit may be in error
when there are no errors in the corresponding data bits. Therefore, the inclusion of
a parity bit with each character would change the probability of receiving a correct
message.
What is the difference between calculating the sample mean and the population mean?
You calculate the actual sample mean, and from that number, you then estimate the probable mean (or the range) of the population from which that sample was drawn.
What is a positively skewed distribution?
A positively skewed or right skewed distribution means that the mean of the data falls to the right of the median. Picturewise, most of the frequency would occur to the left of the graph.
What is the distribution of absolute values of a random normal variable?
It is the so-called "half-normal distribution."
Specifically, let X be a standard normal variate with cumulative distribution function F(z). Then its cumulative distribution function G(z) is given by
Prob(|X| < z)
= Prob(-z < X < z)
= Prob(X < z) - Prob(X < -z)
= F(z) - F(-z).
Its probability distribution function g(z), z >= 0, therefore equals
g(z) = Derivative of (F(z) - F(-z))
= f(z) + f(-z) {by the Chain Rule}
= 2f(z)
because of the symmetry of f with respect to zero. In other words, the probability distribution function is zero for negative values (they cannot be absolute values of anything) and otherwise is exactly twice the distribution of the standard normal.
standard error for proportion is calculated as:
SE = sqrt [(p)(1-p) / n ]
so let us say that "p" is going to represent the decimal proportion of respondents who said YES.... so...
p = 20/25 = 4/5 = 0.8
And... we then are going to say that the complement of "p" which is "1-p" is going to represent the decimal proportion of respondents who said NO ... so... 1-p = 1 - 0.8 = 0.2
Lastly, the "n" in the formula for standard error is equal to 25 because "n" represents the sample size....
So now all you have to do is plug the values you found for "p" and for "1-p"... (remember "p = 0.8" and "1-p = 0.2")... and "n=25"....
Standard Error (SE) = sqrt [(p)(1-p) / n ]
............................ = sqrt [(0.8)(1-0.8) / 25 ]
............................ = sqrt [(0.8)(0.2) / 25 ]
............................ = sqrt [0.16 / 25]
............................ = sqrt (0.0064)
............................ = +/- 0.08
What is the probability of getting a sum of 4 when the two number cubes are rolled?
There are 3 combinations - (1, 3), (2, 2), (3, 1) - that will result in a sum of four. Using fair d6 dice, there are 36 possible combinations - 6 on each die, 6 x 6 = 36. Thus there is a 3/36 = 1/12 = 0.0833.. = 8.3% probability.
What are questions you could ask about probability?
If I have 3 red balls 3green balls 2 white balls & 5black all in one sack what is the probability of finding a white ball?
If a die is rolled 3 times what is probability of having a 5 appear at least twice?
well, there are 216 possibilities, 1/6 x 1/6 x 1/6 = 1/216 so 216 possibilities. and there are 4 ways to get two or more fives, 155, 515, 551, and 555. so 4 out of 216 or 4/216, or 1/54. so there you go.
--- update ---
I disagree. There are 16 ways out of 216 possibilities, so the probability is 16/216 or 2/27. The ways can represent by x55, 5x5, 55x and 555, where x can be any of five numbers not equal to five.
--- actuary fl
If 5 distinct dice are rolled what is the probability of getting at least one 6?
You must notice that it is easier to find the probability that none of the dice are six. This is the complement of the answer. Pc= (55)/(65) = .4019 *the first die has 5 ways to not equal six, and so on for the remaining dice. [55] outcomes have no six. *the total number of outcomes gives each die 6 ways to fall. [65] outcomes overall. *Probability is the favorable outcome divided by the total number of outcomes.
*Probability of an event is 1 minus the complement of the event. P = 1 - .4019 P = .5981 <<<<<<FINAL ANSWER
Why is standard deviation of a statistic called standard error?
The standard error is the standard deviation divided by the square root of the sample size.
P(not raining any day during the week) = 1-4/10 = 3/5 = P(not raining on Sun)*P(not raining on Mon)*P(not raining on Tu)*... = P(not raining on Mon)^7 (assuming all days are equally likely to have rain) so P(not raining on Mon)=(3/5)^(1/7) thus P(raining on Mon) = 1-(3/5)^(1/7)
Generally you should express it as a number between 0 (impossible) and 1 (absolute certainty). Either as a decimal or fraction is OK. For example, the probability of a fair coin toss landing on heads is 0.5. The probability of a fair die landing on 4 is 1/6.
But it's also common to see probability written as a percentage between 0% (impossible) and 100% (absolute certainty). So P(coin landing on heads)=50.0% and P(die landing on 4)=16.7%.
If two dice are thrown what is the probability theat neither one shows a three?
{ (1,1) (1,2) (1,3) (1,4) (1,5) (1,6)
(2,1) (2,2) (2,3) (2,4) (2,5) (2,6)
(3,1) (3,2) (3,3) (3,4) (3,5) (3,6)
(4,1) (4,2) (4,3) (4,4) (4,5) (4,6)
(5,1) (5,2) (5,3) (5,4) (5,5) (5,6)
(6,1) (6,2) (6,3) (6,4) (6,5) (6,6) }
Total number of cases : 36 (from sample space )
Let E be the event where neither of the dice show 3
The number of favorable cases have been striked = 25
p(E) = 25 / 36
To find a real life application of a quadratic function?
When you are trying to find the unknown concentrations in equilibrium reaction ( chemistry ) the result if the ICE table set up devolves into a quadratic equation. Happens in physics to.
Did Blaise Pascal discover the laws of probability?
hi this is Rudd. im bored in power tech.. mr chandlers class.
im sitting here with coltan & Allen. they are talking about how they think they are chubby. which they are not.
but i do know that he did discover the laws of probability.
im doing a stupid power point on it now.
oh, and btw.. coltan thinks Adam Lambert is HOT and you should vote for him.
not like he wont win anyways.
=)
Real life application of a quadratic function?
I think its the dropping of a golf ball off a building! This is because the formula for velocity when something is dropped is a quadratic formula, that is of degree 2.
A random sample is a selection from the population of interest where each item (persons, households, widgets, etc.) has an equal chance of being selected. The idea being that measuring a random sample of sufficient size will accurately (within a margin of error) reflect the "true" value that exists in the population - while at the same time reducing your study to a manageable size.
A random sample is integral in good survey design to reduce bias in your experiment.
Even. as when divided by two it gives a whole number of 14
There are 11+10+17+15+3=56 marbles in total. Of those marbles, 11 are blue and 17 are red, so there are 11+17=28 blue and red marbles. Therefore the probability of choosing a blue or red marble is 28/56=.5, or 50%.
How many different ways can your write the word care?
If you mean in "how many ways can one rearrange the letters in the word care?" (which is what I'll assume as its posted as a maths question) then the answer is 4! ie 4 factorial ie 4x3x2x1=24.